1. A 65 mL of 0.12 M NaF is reacted with 35 mL of 0.12 M Mg(NO3)2. The balanced equation of the reaction is shown below. NaF (aq) + Mg(NO3)2 (aq) → MgF2 (aq) + 2NaNO3 (aq)

a) Write the solubility product constant, Ksp expression for the solubility equilibrium of MgF2. [2 Marks]

b) Determine the molar solubility of MgF2 in the solution. (Ksp for MgF2 = 5.2 x 10-11 ) [8 Marks] c) Calculate the concentration of Mg2+ , F - , NO3 - and Na+ in the final solution. [4 Marks]


1
Expert's answer
2021-07-05T03:28:18-0400

"K_{sp}=[Mg^{2+}][F^-]^2"

At "25^0C" "K_{sp}=5.6\times10^{-11}"


At "275^oC" we should use Van't Hoff equation

"lnK = (\frac{-\Delta H^o}{R})\times \frac{1}{T} + \frac{\Delta S^0}{R}"